### 22 Dec 2023 v1

# Diffusion-Driven Generative Framework for Molecular Conformation Prediction

Abstract
The task of inferring three-dimensional molecular configurations from their two-dimensional graph representations is of critical significance in the domains of computational chemistry and the development of pharmaceuticals. It contributes fundamentally to our grasp of molecular mechanisms and interactions. The rapid evolution of machine learning, especially in the realm of deep generative networks, has catalyzed breakthroughs in the precision of such predictive modeling. Traditional methodologies typically employ a bifurcated strategy: initially estimating interatomic distances followed by sculpting the spatial molecular structure via solving a distance geometry problem. This sequential approach, however, occasionally fails to capture the intricacies of local atomic arrangements accurately, thus compromising the integrity of the resultant structural models. Addressing these deficiencies, this work introduces an avant-garde generative framework: DDGF, which is predicated on the diffusion principles found in classical non-equilibrium thermodynamics. DDGF envisages atoms as discrete entities and is adept at guiding the reversal of diffusion morphing a distribution of stochastic noise back into coherent molecular forms through a process akin to a Markov chain. This transformation begins with the initial representation of a molecular graph in an abstract latent space, progressing to the realization of the three-dimensional forms via an elaborate bilevel optimization scheme, tailored to respect the task’s specific requirements. One of the formidable challenges in this modeling endeavor is the preservation of roto-translational invariance to ensure that the molecular conformations generated respect the laws of physics. Our proposed framework adeptly navigates this complexity, facilitating training from start to finish by fine-tuning a weighted variational lower bound concerning the conformation’s conditional probability. Extensive experiments validate the effectiveness of the proposed DDGF over the state-of-the-art methods.
Index Terms:
Graph Diffusion, Molecular Conformation Prediction, Bilevel Optimization
I Introduction
Graph representation learning has been highly successful in modeling molecules for various tasks, including property prediction [1, 2, 3, 4], time series prediction [5, 6] and molecule generation [7, 8]. In this approach, a molecule is typically represented as an atom-bond graph. Despite its effectiveness in various applications, a more intrinsic and informative representation for molecules is the 3D geometry, also known as conformation, where atoms are represented by their Cartesian coordinates. The 3D structures determine the biological and physical properties of molecules and hence play a key role in many applications such as computational drug and material design [9, 10, 11, 12]. Despite these advances, predicting stable molecular conformation remains a challenging problem. Traditional methods based on molecular dynamics (MD) or Markov chain Monte Carlo (MCMC) are computationally expensive and do not scale well to large molecules [13]. To overcome this challenge, researchers are exploring new methods that combine graph-based models with physics-based simulations or use reinforcement learning to optimize molecular conformations.
Experimentally ascertaining the three-dimensional formations of molecules can be a daunting and resource-intensive process, characterized by both high costs and extended timeframes. Consequently, the quest for computationally deriving accurate and energetically favorable molecular structures has risen to prominence in the realm of computational chemistry. Traditional methodologies in this field have historically depended on techniques such as Markov chain Monte Carlo (MCMC) or molecular dynamics (MD) simulations [14], which involve proposing potential conformations and ascribing to them energy levels via either cost-effective empirical potential models or high-precision quantum chemical computations [15]. The tide has been shifting recently towards incorporating machine learning algorithms to estimate the conditional probability distribution $p(\mathbf{C}|G)$ for generating the stable conformations $\mathbf{C}$ of a given molecular graph $G$, building upon datasets of molecules whose stable conformations are already determined [16, 17]. Notably, two groundbreaking studies [17, 18] have introduced a sequential methodology. The initial phase involves predicting interatomic distances, which sets the stage for the subsequent creation of molecular structures in adherence to these predicted distances, a task that is facilitated by resolving a set of distance geometry challenges[19]. By harnessing the principles of distance geometry, these novel methods manage to capture the essential rotation and translation invariance characteristics of molecular structures, leading to their increasingly accurate predictions of molecular conformations.
The sequential methodologies that first estimate interatomic distances and subsequently deduce conformations encounter a notable drawback: the risk of inaccuracies in local three-dimensional geometries. There are instances where the models for predicting distances might inadvertently lend credence to geometrically incompatible sets of distances, which can cascade into significant deviations. These inaccuracies, when input into the conformation generation stage, can yield atypical and implausible molecular structures. The root of such discrepancies can often be traced back to the distance prediction models, which prioritize the maximization of individual distance likelihoods rather than the holistic accuracy of the resulting conformations. In response to this challenge, our work introduces a streamlined, end-to-end framework that formulates conformation predictions directly from the molecular graph, eliminating the intermediate distance prediction phase. Our approach endeavors to holistically encapsulate the intricate dependencies and interplays among atoms, thus aspiring to elevate the precision and dependability of conformation prediction. This shift towards an integrated predictive model echoes the evolutionary trajectory seen in the domain of protein structure prediction. Innovations exemplified by the AlphaFold2 algorithm underscore the merit and impact of adopting such an end-to-end strategy. AlphaFold2’s refinement over its predecessor, AlphaFold, accentuates the benefits of consolidating distinct phases into a cohesive modeling process [20, 21]. Although AlphaFold2’s internal mechanisms remain partially undisclosed, its profound success in deducing protein structures from linear sequences underscores the potential of end-to-end methodologies in complex structural prediction tasks.
In our research, we introduce DDGF, a cutting-edge, integrated framework tailored for the synthesis of molecular conformations via a sophisticated bilevel programming approach. Central to our method is the integration of roto-translational invariance within the conformational modeling scheme, wherein interatomic distances serve as pivotal variables. Our novel approach transcends the traditional distance prediction models that target minimization of prediction errors in the space of distances. By framing the problem within the context of bilevel optimization [22], DDGF concurrently addresses both the prediction of distances and the resolution of distance geometry challenges that are fundamental to conformational synthesis. To our knowledge, DDGF stands as the inaugural algorithm in the realm of molecular conformation synthesis that is trainable end-to-end while retaining the crucial aspects of roto-translational invariance. Our comprehensive experiments showcase DDGF’s exceptional capabilities, consistently outperforming the current leading methods across a suite of established benchmarks that span conformational synthesis and distance distribution analysis. Furthermore, our findings confirm the criticality of an end-to-end training objective in the creation of conformations that are not only realistic but also chemically significant.
II Related Work
Conformation generation. Recent progress in the field of computational molecular modeling has spawned a plethora of innovative deep generative models and methodologies, each with its unique advantages and inherent challenges. For example, the CVGAE [16] leveraged a variational autoencoder (VAE) for the direct computation of 3D atomic structures. This model, however, did not adequately maintain the critical property of roto-translation equivariance, leading to suboptimal performance. To address this deficiency, newer models have turned to intermediate structural descriptors such as atomic distances and dihedral angles, known for their roto-translationally invariant nature—vital for the precise representation of molecular shapes[23]. Yet, this modular approach, which relies on these intermediate descriptors, introduces certain inefficiencies and complexities during the training phase and generating samples. Approaches like those proposed by [17], [18] involve predicting interatomic distance matrices and subsequently applying Distance Geometry solutions to derive spatial coordinates [19]. Although CONFVAE offered an advancement with its unified, bilevel optimization-based end-to-end framework [24], these techniques invariably encounter the problem of error magnification. The inaccuracies in distance estimations frequently lead to the misguidance of the coordinate determination mechanisms, culminating in the creation of molecular structures that are not just inaccurate but sometimes structurally unfeasible.
To address this problem, CONFGF [25, 26] aimed to learn the gradient of log-likelihood in relation to coordinates. However, in practice, the model still relied on intermediate geometric elements. It estimated the gradient with respect to interatomic distances using denoising score matching (DSM) [27, 28] and then used the chain rule to derive the gradient of coordinates. Unfortunately, by learning the distance gradient via DSM, the model was trained with perturbed distance matrices that might violate the triangular inequality or contain negative values. Consequently, the model learned from invalid distance matrices but was tested with valid ones calculated from coordinates, leading to a serious out-of-distribution problem [29]. A recent work called GEOMOL [30] introduced a highly systematic rule-based pipeline. It focused on predicting a minimal set of geometric quantities, such as lengths and angles, and then reconstructed both local and global structures of conformations through a sophisticated procedure. Additionally, there have been efforts to utilize reinforcement learning for conformation search, exemplified by methods like TorsionNet [31]. However, this approach fundamentally differs from other existing methodologies as it relies on the rigid rotor approximation and only models torsion angles.
Advances in Protein Topology Prediction. The landscape of protein topology prediction has expanded rapidly with a plethora of novel methodologies emerging. Innovations include the use of flow-based models [32], which adeptly map the conformations of protein main chains. Techniques employing recurrent neural architectures have been utilized for sequential amino acid modeling [33].  [34] introduced a novel approach using neural networks to emulate an energy landscape for inferring protein folds. The trailblazing AlphaFold algorithm [20, 35] has notably elevated the benchmark for protein structure prediction accuracy. It is crucial to note, however, that the structural simplicity of proteins, predominantly linear in nature, contrasts starkly with the complex, branched configurations of general molecules that include various ring structures. This structural disparity renders direct application of protein-folding methodologies less suitable for the intricate task of general molecular conformation prediction, underscoring the necessity for distinct modeling techniques that cater to the unique challenges presented by non-protein molecular structures.
Point cloud generation. In the realm of three-dimensional structural synthesis, diffusion-driven models have made their mark, with pioneering efforts seen in recent literature [36, 37]. These approaches have principally targeted the generation of point clouds, a distinct domain from our inquiry. Point clouds typically lack the graph-like structure that embodies diverse atomic and bonding information, a cornerstone of our model’s domain. Furthermore, the principle of equivariance, although not extensively incorporated in these point cloud methodologies, is a cornerstone in our approach. This disparity in foundational concepts underscores the novelty of our model when juxtaposed against the landscape of existing methods.
In summary, the landscape of deep generative models has ushered in a new era for the generation of molecular conformations. Each innovation offers unique benefits and faces distinct challenges. There is a continuous need for exploration to surmount the persisting hurdles in enhancing the precision and computational efficiency of these conformational generation techniques.
III Preliminaries
III-A Notations and Problem Definition
Notations. In this paper, we delineate a molecule comprising $n$ atoms as a graph $G=(V,E)$ without directionality. Here, $V={v_{i}}{i=1}^{n}$ signifies the vertices corresponding to atoms, and $E={e{ij}|(i,j)\in{1,\ldots,n}^{2}}$ represents the bonds between them as edges. Every vertex $v_{i}$ encapsulates atomic properties such as atomic number or element category. Interatomic relationships are denoted by edges $e_{ij}$, each annotated with its bond type. Edges without a chemical bond are designated a distinct virtual category. Geometrically, vertices in $V$ are each associated with a 3D coordinate vector $\mathbf{c}_{i}\in\mathbb{R}^{3}$, situating the atom in a tridimensional framework. The entirety of these coordinates forms the molecular conformation matrix $\mathbf{C}=[\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{n}]$ within $\mathbb{R}^{n\times 3}$ space.
Problem Definition. The task of molecular conformation generation is a conditional generative problem, where we are interested in generating stable conformations for a provided graph $G$. Consider a set of molecular graphs $G$ with each graph’s conformation $\mathbf{C}$ acting as independently and identically distributed samples emanating from a latent Boltzmann distribution [32]. The crux of our endeavor lies in devising a generative model denoted by $p_{\theta}(\mathbf{C}|G)$. This model, which should facilitate straightforward sample generation, is tasked with emulating the probabilistic nature of the Boltzmann distribution.
III-B Equivariance
Equivariance plays a vital role in machine learning models that deal with atomic systems. It is the principle that certain properties of atoms, such as the vectors representing atomic dipoles or forces, should transform consistently with the transformations applied to the atomic coordinates. For example, if a molecule is rotated in space, the direction of the dipoles or forces should rotate in the same way [9, 38, 39, 40, 12]. Incorporating this inductive bias directly into the architecture of a machine learning model has been proven to be highly effective for accurately modeling three-dimensional geometries. By ensuring that the model’s outputs change predictably with changes to the inputs, we can greatly enhance the model’s ability to generalize from one situation to another [23]. This is particularly important in the field of molecular modeling, where the ability to predict how a molecule will behave under various spatial transformations is crucial. Formally, a function F : X → Y is equivariant w.r.t a group $G$ if:
$\mathcal{F}\odot T_{g}(x)=S_{g}\odot\mathcal{F}(x),$ (1)
where $T_{g}$ and $S_{g}$ are transformations for an element $g\in G$, acting on the vector spaces $X$ and $Y$, respectively.
III-C Equivariant Graph Neural Networks
EGNNs [41] are a type of Graph Neural Network that satisfies the equivariance constraint. In this work, we consider interactions between all atoms, and therefore assume a fully connected graph $G$ with nodes $v_{i}\in V$. Each node $v_{i}$ is endowed with coordinates $\mathbf{c}_{i}\in\mathbb{R}^{3}$ as well as features $h_{i}\in\mathbb{R}^{d}$. In this setting, EGNN consists of the composition of Equivariant Convolutional Layers $\mathbf{c}^{l+1},\mathbf{h}^{l+1}=EGCL[\mathbf{c}^{l},\mathbf{h}^{l}]$ which are defined as:
$\mathbf{m}_{ij}=\phi_{e}$ $(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},d_{ij}^{2},a_{ij}),\quad\mathbf{h}_{i}^{l+1}=\phi_{h}(\mathbf{h}_{i}^{l},\sum_{j\neq i}\tilde{e}_{ij}\mathbf{m}_{ij}),$ (2)
$\mathbf{c}_{i}^{l+1}$ $=\mathbf{c}_{i}^{l}+\sum_{j\neq i}\frac{\mathbf{c}_{i}^{l}-\mathbf{c}_{j}^{l}}{d_{ij}+1}\phi_{x}(\mathbf{h}_{i}^{l},\mathbf{h}_{j}^{l},d_{ij}^{2},a_{ij}),$
where $l$ indexes the layer, and $d_{ij}=||\mathbf{c}^{l}_{i}-\mathbf{c}^{l}_{j}||_{2}$ is the euclidean distance between nodes $(v_{i},v_{j})$, and $a_{ij}$ are optional edge attributes. The difference $(\mathbf{c}^{l}_{i}-\mathbf{c}^{l}_{j})$ in Equation 2 is normalized by $d_{ij}+1$ as done in [42] for improved stability, as well as the attention mechanism which infers a soft estimation of the edges $\tilde{e}_{ij}=\phi_{inf}(\mathbf{m}_{ij})$. All learnable components ($\phi_{e}$,$\phi_{h}$, $\phi_{x}$ and $\phi_{inf}$ ) are parametrized by fully connected neural networks. An entire EGNN architecture is then composed of $L$ EGCL layers which applies the following non-linear transformation $\hat{\mathbf{c}},\hat{\mathbf{h}}=EGNN[\mathbf{c}^{0},\mathbf{h}^{0}]$.
III-D Bilevel Optimization
Bilevel programs are defined as optimization problems where a set of variables involved in the (outer) objective function are obtained by solving another (inner) optimization problem. Formally, given the outer objective function F and the inner objective H, and the corresponding outer and inner variables $\theta$ and $w$, a bilevel program can be formulated by:
$\min_{\theta}F(w_{\theta})\ such\ that\ w_{\theta}\in arg\min_{w}H(w,\theta).$ (3)
Bilevel programs have shown effectiveness in a variety of situations such as hyperparameter optimization, adversarial and multi-task learning, as well as meta-learning.
Typically solving equation 3 is intractable since the solution sets of $w_{\theta}$ may not be available in closed form. A common approach is to replace the exact minimizer of the inner object $H$ with an approximation solution, which can be obtained through an iterative optimization dynamics $\Phi$ such as stochastic gradient descent (SGD). Starting from the initial parameter $w_{0}$, we can get the approximate solution $w_{\theta,T}$ by running $T$ iterations of the inner optimization dynamics $\Phi$, i.e., $w_{\theta,T}=\Phi(w_{\theta,T-1},\theta)=\Phi(\Phi(w_{\theta,T-2},\theta),\theta)$, and so on. In the general case where $\theta$ and $w$ are real-valued and the objectives and optimization dynamics is smooth, the gradient of the object $F(w_{\theta,T})$ w.r.t. $\theta$, named hypergradient $\nabla_{\theta}F(w_{\theta,T})$, can be computed by:
$\nabla_{\theta}F(w_{\theta,T})=\partial_{w}F(w_{\theta,T})\nabla_{\theta}w_{\theta,T},$ (4)
where $\partial$ denotes the partial derivative to compute the Jacobian on immediate variables while $\nabla$ denotes a total derivative taking into account the recursive calls to $F$. The above gradient can be efficiently calculated by unrolling the optimization dynamics with back-propagation, i.e., reverse mode automatic differentiation  [43], where we repeatedly substitute $w_{\Phi,t}=\Phi(w_{\theta,t-1},\theta)$ and apply the chain rule.
IV Proposed Method
Figure 1: The overall framework of DDGF.
In our framework, the authentic conformation $\mathbf{c}^{0}$ serves as a reference for a series of transformations through latent space, defined by $\mathbf{c}^{t}$ for $t=1,\cdots,T$, where $t$ delineates the progression of diffusion steps. This is governed by a dual-phase probabilistic model [44]. Here, the initial phase disperses incremental noise through the conformational states, effectively transitioning from the true state to a noise-saturated latent space. Counteracting this, the subsequent phase engages in a systematic noise abatement, iteratively refining these perturbed states towards reconstructing the pristine conformation $\mathbf{c}^{0}$. A conceptual diagram outlining these mechanisms is depicted in Figure1.
IV-A Diffusion Process
Drawing parallels from thermodynamics, we envision the molecules’ conformation C as particles within an ever-fluctuating system. As moments elapse, the pristine equilibrium state $\textbf{c}^{0}$ embarks on a stochastic journey through a series of increasingly disordered states $\textbf{c}^{t}$, culminating in a state of maximal entropy akin to white noise after $T$ transitional steps. This progression, distinctive from conventional latent variable constructs, is characterized in our diffusion model by a predetermined posterior distribution $q(\mathbf{c}^{1:T}|\mathbf{c}^{0})$, which remains constant throughout the learning process. Specifically, we define it as a Markov chain according to a fixed variance schedule $\beta_{1},\cdots,\beta_{T}$ :
$q(\mathbf{c}^{1:T}|\mathbf{c}^{0})$ $=\Pi_{t=1}^{T}q(\mathbf{c}^{t}|\mathbf{c}^{t-1}),$ (5)
$q(\mathbf{c}^{t}|\mathbf{c}^{t-1})=\mathcal{N}$ $(\mathbf{c}^{t};\sqrt{1-\beta_{t}}\mathbf{c}^{t-1},\beta_{t}I).$
Note that, in this work we do not impose specific (invariance) requirement upon the diffusion process, as long as it can efficiently draw noisy samples for training the generative process $p_{\theta}(\mathbf{c}^{0})$.
Defining $\alpha_{t}=1-\beta_{t}$, and ${\bar{\alpha}}_{t}=\Pi^{t}_{s=1}\alpha_{s}$, we observe a noteworthy aspect of the forward process: the conditional distribution $q(\mathbf{c}^{t}|\mathbf{c}^{0})$ at any given $t$ can be precisely expressed as $q(\mathbf{c}^{t}|\mathbf{c}^{0})=\mathcal{N}(\mathbf{c}^{t};\sqrt{{\bar{\alpha}}_{t}}\mathbf{c}^{0},(1-{\bar{\alpha}}_{t})I)$. With an adequately extended series up to $T$ steps, this process inherently transforms $\mathbf{c}^{0}$ into a standard Gaussian distribution, implying a harmonious choice for $p(\mathbf{c}^{T}$ would be an isotropic Gaussian distribution.
IV-B Reverse Process
Our objective is to infer the original molecular conformations, denoted by $\mathbf{c}^{0}$, starting from a state of maximum entropy represented by white noise $\mathbf{c}^{T}$, conditioned on the provided molecular graph structures $G$. We approach this inference task as the temporal reversal of the diffusion sequence, initiating from the high-entropy state where the conformational coordinates $\mathbf{c}^{T}$ are distributed according to $p(\mathbf{c}^{T})$. We formulate this reverse dynamics as a conditional Markov chain with learnable transitions:
$p_{\theta}(\mathbf{c}^{0:T-1}|G,\mathbf{c}^{T})=\Pi_{t=1}^{T}p_{\theta}(\mathbf{c}^{t-1}|G,\mathbf{c}^{t}),$ (6)
$p_{\theta}$ $(\mathbf{c}^{t-1}|G,\mathbf{c}^{t})=\mathcal{N}(\mathbf{c}^{t-1};\mu_{\theta}(G,\mathbf{c}^{t},t),\sigma_{t}^{2}I).$
In our framework, neural networks parameterized by $\theta$ are employed to approximate mean values $\mu_{\theta}$, while $\sigma_{t}$ represents a variance that can be arbitrarily chosen by the practitioner. The process commences with the assignment of the initial state distribution $p(\mathbf{c}^{T})$ as a standard Gaussian distribution. For any given molecular graph $G$, its three-dimensional configuration is constructed by first sampling disordered state particles $\mathbf{c}^{T}$ from $p(\mathbf{c}^{T})$, which are subsequently progressively sculpted into the desired structure via a sequence of reverse Markov transitions $p_{\theta}(\mathbf{c}^{t-1}|G,\mathbf{c}^{t})$.
[…]
IV-C Training Objective
In order to account for the multiplicity of stable molecular structures, our study implements a conditional variational autoencoder (CVAE) framework to construct the probabilistic model $p(\mathbf{C}|G)$. Within this framework, we introduce a latent variable $z$ to encapsulate the intrinsic uncertainties associated with the generation of molecular conformations. The architecture of the CVAE comprises a prior latent distribution $p_{\psi}(z|G)$ along with a decoding mechanism $p_{\theta}(\mathbf{C}|z,G)$ which delineates the conditional likelihood of the conformation $\mathbf{C}$ for a given $z$. Concurrently, the training phase is augmented with an inferential component (encoder) $q_{\psi}(z|\mathbf{C},G)$, enhancing the model’s ability to learn from data. The encoder and decoder are jointly trained to maximize the evidence lower bound (ELBO) of the data log-likelihood:
$\mathbb{E}$ $[logp_{\theta}(\mathbf{c}^{0}|G)]\geq\mathbb{E}\left[log\mathbb{E}_{q(\mathbf{c}^{1:T}|\mathbf{c}^{0})}\frac{p_{\theta}(\mathbf{c}^{0:T}|G)}{q(\mathbf{c}^{1:T}|\mathbf{c}^{0})}\right]$ (7)
$\quad-D_{KL}[q_{\phi}(z|\mathbf{C},G)||p_{\psi}(z|G)]$
$:=-\mathcal{L}_{ELBO}.$
The ELBO, fundamental to our model’s training regime, is essentially the sum of two distinct components: the negative reconstruction loss, $L_{recon}$, and the latent space regularization term, $L_{prior}$. The former measures the fidelity of reconstructed data to the original, while the latter ensures the latent variable distribution remains well-behaved, or close to the prior. To encapsulate the model’s probabilistic nature, both the encoder’s output $q_{\phi}(z|\mathbf{C},G)$ and the prior distribution $p_{\psi}(z|G)$ are formulated as diagonal Gaussian distributions $\mathcal{N}(z|\mu_{\phi}(\mathbf{C},G),\sigma_{\phi}(\mathbf{C},G))$ and $\mathcal{N}(z|\mu_{\psi}(G),\sigma_{\psi}(G))$, respectively. The parameters of these distributions, namely the means and standard deviations, are deduced by specialized graph neural networks. In the process of optimizing the ELBO, a reparametrization trick is employed to enable backpropagation through the stochastic sampling process. Specifically, a latent variable $z$ is sampled by first calculating $\mu_{\phi}(\mathbf{C},G)$ and $\sigma_{\phi}(\mathbf{C},G)$, then adding a scaled random noise component $\epsilon$, where $\epsilon$ follows a standard Gaussian distribution $\mathcal{N}(0,I)$. This methodology facilitates gradient-based optimization by making the sampling process differentiable.
IV-D Inner Objective
To circumvent the rotational and translational dependency inherent in Cartesian coordinate generation for molecular conformations, modern methods [17] typically employ an alternate strategy. These methods utilize a decoder architecture designed to produce inter-atomic distance matrices, with $d_{\theta,\phi}=D_{\theta}(z_{\phi},G)$ representing the decoded distances conditioned on the latent variable $z$ and the graph $G$. These distance matrices serve as a translationally and rotationally invariant representation of the molecule’s geometry. After the distance matrix $d$ is generated, it can then be used as a basis for reconstructing the three-dimensional conformation $\mathbf{C}$. This reconstruction is achieved by addressing a distance geometry problem, which essentially entails finding a set of coordinates that best comply with the given inter-atomic distances. The optimization process involved is often a complex numerical problem that requires sophisticated techniques to resolve the coordinates in a way that faithfully respects the distances while also conforming to physical plausibility.
$R_{\theta,\phi}$ $=arg\min_{\mathbf{C}}H(\mathbf{C},D_{\theta}(z_{\phi},G))$ (8)
$=arg\min_{\mathbf{C}}H(\mathbf{C},d_{\theta,\phi})$
$=arg\min_{\mathbf{C}}\left[\sum_{e_{uv}\in E}(||r_{u}-r_{v}||_{2}-d_{uv})^{2}\right].$
IV-E Outer Objective
Our primary aim is to refine the generation of 3D molecular structures, aligning them as closely as possible to the actual models, regardless of their orientation or positioning. To measure this alignment’s precision, we employ the post-alignment Root-Mean-Square Deviation (RMSD), a standard gauge in the field. This metric is determined after optimally aligning a generated conformation $\mathbf{C}$ to a reference $\mathbf{C}^{*}$ using an alignment operation $A$, ensuring minimal disparity between the two. The RMSD is thus defined as the square root of the average squared distance between the atoms of the aligned generated conformation and the reference after this operation is applied.
$RMSD(\mathbf{C},\hat{\mathbf{C}})=(\frac{1}{n}\sum_{i=1}^{n}||\mathbf{c}_{i}-\hat{\mathbf{c}}_{i}||^{2})^{\frac{1}{2}}.$ (9)
where $n$ is the number of atoms. Then the reconstruction objective $L_{recon}$ can be written as:
$F(R_{\theta,\phi})$ $=logp_{\theta}(\mathbf{C}|z,G)$ (10)
$=-\sum_{i=1}^{n}\sum_{j=1}^{3}(\mathbf{C}_{ij}-A(\mathbf{C},\mathbf{C}^{*})_{ij})^{2},$
which is the outer loop objective for computing the reconstruction loss and maximize the log-likelihood.
IV-F Bilevel Program
To navigate the complexities of molecular conformation generation, our approach adopts a nested optimization scheme, where we distinguish between two intertwined objectives: one at a macro-level and another at a micro-level. The macro-level objective (denoted as the outer problem in Equation 10) is to capture the conditional distribution $p(\mathbf{C}|G)$ of molecular structures, encapsulating the full breadth of conformations that molecules can adopt given their underlying graph $G$. Concurrently, the micro-level objective (referred to as the inner problem in Equation 8) is dedicated to determining the precise molecular conformation from a collection of interatomic distances that are predicted. By taking the expectation over latent variable $z$, the resulting bilevel program for calculating the reconstruction term $L_{reconin}$ equation 7 can be written as:
$\max_{\theta,\phi}\mathbb{E}_{z\sim q_{\phi}(z|\mathbf{C},G)}[F(\mathbf{C}_{\theta,\phi},\theta)]$ (11)
$such\ that\ \mathbf{C}_{\theta,\phi}=arg\min_{\mathbf{C}}H(\mathbf{C},D_{\theta}(z_{\phi},G)).$ (12)
Addressing the complexity of the bilevel optimization in our model, we circumvent the lack of closed-form solutions and intractable expectations by leveraging a variational inference approach. This is paired with the reparametrization trick to provide a tractable, gradient-based estimation of molecular conformations, a technique we detail in the subsequent sections.
IV-G Generative Model
Our generative model for molecular conformations is underpinned by the use of message-passing neural networks (MPNNs) [45], a class of graph neural networks renowned for their exemplary representation learning capabilities for molecular structures. MPNNs work on the principle of updating atom embeddings through an aggregation of neighboring node information within the graph structure $G$, ensuring invariance to graph isomorphism and facilitating state-of-the-art molecular modeling [46, 47, 48]. We have adopted the MPNN architecture akin to the one delineated [17] for the encoder $q_{\phi}(z|\mathbf{C},G)$ and the prior $p_{\psi}(z|G)$. Recognizing the intensive memory demands of bilevel optimization, our decoder $p_{\theta}(\mathbf{C}|z,G)$ is innovatively implemented using a continuous normalizing flow (CNF) based on an ordinary differential equation (ODE), offering the advantage of a fixed memory footprint regardless of the number of layers. Details of our decoder’s design are provided in the subsequent sections.
In our design, the decoder operates through two sequential stages. The initial stage involves a distance prediction network $D_{\theta}(z,G)$ which interprets the latent space variable $z$ into an array of interatomic distances, denoted by $d$. Subsequently, the model employs a differential geometry algorithm that is capable of reconstructing the molecular conformation $\mathbf{C}$ from the set of distances $d$. The network $D_{\theta}(z,G)$ is a conditional adaptation of the CNF model which commences with a noise distribution $d(t_{0})$ representing the starting distances in the CNF ODE path originating from a standard Gaussian distribution $\mathcal{N}(0,I)$ and culminates with the ultimate distances $d=d(t_{1})$. The transformation is conditioned on the latent variable $z$ as well as the graph $G$:
$d$ $=D_{\theta}(z,G)$ (13)
[…]
$\mathcal{L}_{aux}$ $=logp_{\theta}(d|z,G)$ (14)
[…]
In the training of our molecular conformation generative model, we employ an ODE solver to integrate the continuous-time dynamics for backpropagation, complemented by an auxiliary objective $L_{aux}$ that imposes a regularizing effect on distance predictions. This integration ensures that the model’s parameters are optimized to yield accurate molecular geometries, harnessing the ODE solver’s gradient estimations and the guiding influence of $L_{aux}$ to refine the distance predictions crucial for generating realistic molecular conformations. In summary, the training objective can be interpreted as the sum of three parts:
$\mathcal{L}(\theta,\phi,\psi)=\mathcal{L}_{recon}+\lambda\mathcal{L}_{prior}+\alpha\mathcal{L}_{aux},$ (15)
where $\lambda$ and $\alpha$ are hyperparameters.
IV-H Bilevel Optimization
To tackle the optimization of the bilevel problem articulated in equation 11 and equation 12, we propose a pragmatic algorithm that builds on existing methods for inferring 3D structures from interatomic distances. The inner problem, represented by equation 12, is akin to traditional distance geometry challenges commonly encountered in molecular structure determination. Analogous problems have been approached through semi-definite programming (SDP) for deducing protein configurations from NMR data, as well as employing an Alternating Direction Method of Multipliers (ADMM) to computationally fold proteins into their native three-dimensional forms. Drawing inspiration from these methodologies, our algorithm is designed to effectively navigate the complexity of the inner optimization problem, thereby facilitating accurate reconstruction of molecular geometries from predicted interatomic distances. This approach allows us to efficiently generate 3D coordinates that are consistent with the given pairwise distances, which is critical for the successful application of our model to the prediction and generation of molecular conformations.
In this initial work we choose gradient descent (GD), with tractable learning dynamics $\Phi$, to approximately solve for the geometry:
$\mathbf{C}_{\theta,\phi,t+1}=\Phi(\mathbf{C}_{\theta,\phi,t},d_{\theta,\phi})=\mathbf{C}_{\theta,\phi,t}-\eta\nabla H(\mathbf{C}_{\theta,\phi,t},d_{\theta,\phi}),$ (16)
where $\eta$ is the learning rate and $d_{\theta,\phi}$ is the distance set generated from the distance prediction model. Under appropriate assumptions and for a number of updates $t\to\infty$, GD can converge to a proper geometry $\mathbf{C}_{\theta,\phi}$ that depends on the predicted pairwise distances.
In the process of model training, we concentrate on determining the hypergradient by computing the derivative of the expected reconstruction error with respect to the parameters, which influences the adjustment of the model to enhance conformation predictions. This is achieved by generating molecular conformations using a series of gradient descent steps, signified by $\mathbf{C}_{\theta,\phi,T}$, to approximate the solution for the distance geometry problem.
Now we can write the hypergradient as:
$\nabla_{\theta,\phi}$ $\mathbb{E}_{z\sim q_{\phi}(z|\mathbf{C},G)}[F(\mathbf{C}_{\theta,\phi,T})]$ (17)
$=\mathbb{E}_{z\sim q_{\phi}(z|\mathbf{C},G)}\partial_{\mathbf{C}}[F(\mathbf{C}_{\theta,\phi,T})]\nabla_{\theta,\phi}\mathbf{C}_{\theta,\phi,T}.$
To compute the gradient $\nabla_{\theta,\phi}\mathbf{C}_{\theta,\phi,T}$, we fully trace the evolution of the inner optimization from the end state $\mathbf{c}_{T}$ back to the initial state $\mathbf{c}_{0}$. During the forward pass, we store each intermediary geometry $\mathbf{c}_{0,\cdots,T}$ resulting from the inner optimization dynamics. For the backward computation, these saved geometries facilitate the calculation of gradients via a sequence of Vector-Jacobian Products (VJPs). During the reverse computation, the gradient starting from the $\partial_{\mathbf{c}_{T}}F$ can be propagated to the intermediate geometries $\mathbf{c}_{t}$ through $\nabla_{\mathbf{c}_{t}}\mathbf{c}_{t+1}$:
$\nabla_{\mathbf{c}^{t}}\mathbf{c}^{t+1}$ $=\nabla_{\mathbf{c}^{t}}(\mathbf{c}^{t}-\eta\nabla_{\mathbf{c}^{t}}H(d_{\theta,\phi},\mathbf{c}^{t}))$ (18)
$=1-\eta\nabla_{\mathbf{c}^{t}}^{2}H(d_{\theta,\phi},\mathbf{c}^{t}).$
The notation $\nabla_{\mathbf{c}^{t}}^{2}$ represents the Hessian matrix with respect to the conformation $\mathbf{c}^{t}$. By iteratively calculating the derivatives $\nabla_{\mathbf{c}^{t}}\mathbf{c}^{T}$, we can determine the adjoint values for the distances $d_{\theta,\phi}$. These adjoints are then used to perform Vector-Jacobian Products (VJPs), allowing the gradients to be propagated back through the network, affecting the parameters of the encoder $q_{\phi}$and the decoder $p_{\theta}$. Formally, $\nabla_{d}\mathbf{c}^{T}$ is computed by:
$\nabla_{d_{\theta,\psi}}\mathbf{c}^{T}$ $=\sum_{t=T-1}^{0}[\nabla_{\mathbf{c}^{t+1}}\mathbf{c}^{T}]\nabla_{d}\mathbf{c}^{t+1}$ (19)
$=-\eta\sum_{t=T-1}^{0}[\nabla_{\mathbf{c}^{t+1}}\mathbf{c}^{T}]\nabla_{d}(\nabla_{\mathbf{c}^{t}}H(d_{\theta,\phi},\mathbf{c}^{t})),$
where $\nabla_{\mathbf{c}^{t+1}}\mathbf{c}^{T}$ can be substituted by equation 18. The computation can be done efficiently with reverse-mode automatic differentiation software such as PyTorch.
IV-I Sampling
To synthesize a molecular conformation $\mathbf{C}$ from a given molecular graph $G$, we first sample a latent variable $\tilde{z}$ from the learned prior distribution $p_{\psi}(z|G)$. Next, we obtain a set of initial interatomic distances $d(t_{0})$ by sampling from a Gaussian distribution. These sampled distances $\tilde{d}(t_{0})$ are processed by the trained Neural ODE $G_{\theta}$, which, given the conditions $\tilde{z}$ and $G$, transforms $\tilde{d}(t_{0})$ into a new set of distances $\tilde{d}$. Finally, we use these predicted distances $\tilde{d}$ to reconstruct the molecular conformation $\mathbf{C}$ by solving the optimization problem $\text{arg min}_{\mathbf{C}}H(\mathbf{C},d_{\theta,\phi})$, which aims to find the conformation $R$ that best matches the distances, as defined in the previously referred equation.
Algorithm 1 Training Algorithm of DDGF
0:  objective reweighting coefficients $\alpha$ and $\lambda$; the inner loop optimization iterations $T$ and learning rate $\eta$; alignment function $A(\textperiodcentered,\textperiodcentered)$; data samples $\{G_{t},\mathbf{c}^{*}_{t}\}$.
0:  prior $p_{\psi}(z|G)$, decoder $p_{\theta}(\mathbf{C}|z,G)$ and its dynamics defined as $g_{\theta}$ , encoder $q_{\phi}(z|\mathbf{C},G)$.
1:  Sample $\mathbf{c}^{T}\sim p(\mathbf{c}^{T})=\mathcal{N}(0,I)$
2:  for $s=T,T-1,\cdots,1$ do
3:     Compute $\mu_{\theta}(\mathbf{c}^{s},G,s)$ from $\epsilon_{\theta}(\mathbf{c}^{s},G,s)$;
4:     Sample $\mathbf{c}^{s-1}\sim\mathcal{N}(\mathbf{c}^{s-1};\mu_{\theta}(\mathbf{c}^{s},G,s),\sigma^{2}_{t}I)$
5:     Calculate $\mu_{q},\sigma_{q}$;
6:     Calculate $\mathcal{L}_{prior}$;
7:     $d^{*}\leftarrow R_{t}^{*}$;
8:     Calculate $\mathcal{L}_{aux}$;
9:     Initialize $\mathbf{0}^{0}$, sample $d(t_{0})\sim\mathcal{N}(0,I)$;
[…]
11:     for $t=1,2,\cdots,T$ do
12:        $\mathbf{c}^{t+1}=\mathbf{c}^{t}-\eta\nabla H(\mathbf{c}^{t},d)$;
13:     end for
14:     $R\leftarrow\mathbf{c}^{T}$;
15:     Calculate $\mathcal{L}_{recon}$;
16:     Calculate $\mathcal{L}$;
17:     Update $\theta,\phi,\psi$;
18:  end for
19:  return  $q_{\phi},p_{\theta},p_{\psi}$
V Experiments
In the ensuing paragraphs, we subject DDGF to rigorous empirical testing, focusing on its efficacy in generating equilibrium conformations for molecules of various sizes, including those that are pertinent to pharmacological interests. In line with established precedents in the field [25, 30], we scrutinize our approach alongside other prominent contenders across two benchmark tests. The first, Conformation Generation, gauges the proficiency of the model in capturing the conformational distribution by assessing both the diversity and the fidelity of the conformations it generates. The second, Property Prediction, examines the model’s ability to infer molecular properties accurately. We outline the general setup for our experiments before delving into the evaluation criteria tailored to each task and the subsequent discussion of the findings.
V-A Experimental Settings
V-A1 Datasets
Building on the frameworks set by recent investigations [18], our analysis incorporates data from the GEOM-QM9 [49] and GEOM-Drugs [30] datasets, each hosting a spectrum of molecular structures. GEOM-QM9 is renowned for its collection of small molecules, whereas GEOM-Drugs encompasses medium-sized organic compounds. We adopt the dataset partitions established by Shi et al. [25], with both datasets’ training splits encompassing 40,000 molecules, each represented by 5 distinct conformations, cumulating in a total of 200,000 conformations. The validation set mirrors the size of the training split. For our evaluation, the test split presents 200 unique molecules, translating to 22,408 conformations within QM9 and 14,324 for the Drugs dataset. This rich repository of structural data sets the stage for our comprehensive model assessments.
The GEOM-QM9 dataset extends the original QM9 collection by including multiple conformations for each molecule, focusing on smaller compounds with up to 9 heavy atoms. For our training, we extracted 50,000 pairs of conformations and molecules from this dataset, and set aside 17,813 conformations from 150 molecular structures for testing. In contrast, the GEOM-Drugs dataset comprises larger molecules with an average of 6.5 rotatable bonds, indicative of complex molecular dynamics. From this, we formed our training set with another 50,000 pairs and allocated 9,161 conformations from 100 molecular structures for testing, allowing us to assess our model’s performance across a wide range of molecular complexities.
V-A2 Evaluation metrics
Evaluating the conformation generation models hinges on their ability to produce diverse and accurate molecular structures, for which the root-mean-square deviation (RMSD) serves as a critical measure after aligning the structures using the Kabsch algorithm. Metrics include Coverage, which reflects the model’s diversity by indicating the proportion of reference conformations matched within a certain RMSD threshold, and Matching, which mirrors precision by showing how many generated conformations closely resemble those in the reference set. Together with the Minimum and Maximum RMSD metrics that respectively shed light on the average best-case alignment and the worst outliers, these tools collectively gauge the model’s efficacy in crafting conformations that are both varied and true to real molecular shapes.
$COV-R(S_{g},S_{r})=\frac{1}{|S_{r}|}\left|\{\mathbf{c}\in S_{r}|RMSD(\mathbf{c},\hat{\mathbf{c}})\leq\delta,\hat{\mathbf{c}}\in S_{g}\}\right|,$
$MAT-R(S_{g},S_{r})=\frac{1}{|S_{r}|}\sum_{\mathbf{c}\in S_{r}}\min_{\hat{\mathbf{c}}\in S_{g}}RMSD(\mathbf{c},\hat{\mathbf{c}}),$ (20)
[…]
The property prediction task appraises the molecular ensemble properties over a range of generated conformations, providing a direct evaluation of sample quality. Specifically, we adopt the approach detailed in previous research where a subset of 30 molecules from the GEOM-QM9 dataset is used to generate 50 samples for each molecule. Utilizing the chemical computation toolkit PSI4, we calculate the energy ($E$) and HOMO-LUMO gap ($\epsilon$) for each conformer. These calculations allow us to compare several statistics, the average energy ($\bar{E}$), the lowest energy conformer ($E_{min}$), the average HOMO-LUMO gap ($\bar{\Delta\epsilon}$), the smallest gap ($Delta\epsilon_{min}$), and the largest gap ($\Delta\epsilon_{max}$), against the ground truth values. Discrepancies in these values serve as indicators of the generative model’s accuracy in reflecting true molecular properties.
V-B Baselines
The evaluation of our method, denoted as DDGF, is against four recent or established state-of-the-art baselines, inclusive of machine learning approaches and classic computational methods:
  • CVGAE [16]: This is a conditional Variational Autoencoder (VAE)-based model that employs several layers of graph neural networks to encode the atom representation from a molecular graph. It then directly predicts the 3D coordinates of molecules.
  • GRAPHDG [17]: Also utilizing a conditional VAE framework, this model differs from CVGAE in that it does not generate conformations directly in 3D space. Instead, it learns the distribution over interatomic distances and then employs a distance geometry algorithm to convert these distances into molecular conformations.
  • CGCF [18]: Standing for Continuous Geometric Conformational Flow, this two-stage method predicts atomic pairwise distances using continuous normalizing flows. The distances are then utilized to generate the conformations.
  • RDKIT [50]: A classical and widely utilized approach to conformation generation, RDKIT is an open-source toolkit that leverages a traditional distance geometry method with a comprehensive calculation of edge lengths derived from computational chemistry.
These baseline methods represent a spectrum from classical computational chemistry to modern machine learning models, providing a robust comparison across different strategies for molecular conformation generation. DDGF is tested against these baselines to determine its effectiveness and accuracy in generating realistic and diverse molecular conformations.
V-C Results and Analysis
TABLE I: Comparison of different methods on the conformation generation task.
Dataset GEOM-QM9 GEOM-Drugs
[…]
Mean Median Mean Median Mean Median Mean Median
CVGAE 8.52 5.62 0.7810 0.7811 0.00 0.00 2.5225 2.4680
GraphDG 55.09 56.47 0.4649 0.4298 7.76 0.00 1.9840 2.0108
CGCF 69.60 70.64 0.3915 0.3986 49.92 41.07 1.2698 1.3064
ConfVAE 75.57 80.76 0.3873 0.3850 51.24 46.36 1.2487 1.2609
DDGF 77.33 82.37 0.3742 0.3729 54.47 49.81 1.0254 1.0344
RDKit 79.94 87.20 0.3238 0.3195 65.43 70.00 1.0962 1.0877
CVGAE+FF 63.10 60.95 0.3939 0.4297 83.08 95.21 0.9829 0.9177
GraphDG+FF 70.67 70.82 0.4168 0.3609 84.68 93.94 0.9129 0.9090
CGCF+FF 73.52 72.75 0.3131 0.3251 92.28 98.15 0.7740 0.7338
ConfVAE+FF 77.95 79.14 0.2851 0.2817 91.48 99.21 0.7743 0.7436
DDGF+FF 78.42 81.67 0.2745 0.2718 92.47 99.22 0.7716 0.7219
The summarized results in Table I depict a comprehensive performance assessment of molecular conformation generation methods. The table is bifurcated into two sections, where the top five rows feature deep generative models, and the bottom six rows include various methods enhanced with an additional rule-based force field for optimizing the generated structures. The reported metrics include Coverage (COV) and Matching (MAT) scores, with Mean and Median values calculated across different molecular graphs within the GEOM test set. For a fair and consistent evaluation [18], the size of the generated set is double that of the reference set. The standout performance of DDGF, as reflected in the tables, indicates its exceptional capability to model the multimodal distribution of molecular conformations, delivering both precision and diversity. This is especially evident with the more complex and larger molecules found in the Drugs dataset, where DDGF surpasses state-of-the-art machine learning (ML) models by a significant margin. An interesting observation is the slight edge of DDGF-C over DDGF-A, hinting at the efficacy of the chain-rule approach in enhancing the optimization process. Given this finding, DDGF-C is selected for further comparisons.
On the intricate Drugs dataset, the performance of RDKIT was also evaluated. According to the data presented in Table I, there is a consensus with previous studies that state-of-the-art machine learning (ML) models surpass RDKIT in Coverage-Recall (COV-R) and Matching-Recall (MAT-R) metrics. Nevertheless, when it comes to the newly introduced Precision-based metrics, ML models still fall short. This discrepancy suggests that while ML models are more exploratory in identifying a broader range of potential structures, RDKIT tends to focus on producing a smaller set of the most probable structures, emphasizing quality over quantity. The tendency of RDKIT to produce higher quality at the expense of diversity has been previously attributed to its use of an empirical force field (FF), a strategy employed in traditional computational chemistry to refine and optimize molecular structures. In an attempt to level the playing field and conduct a more equitable comparison, DDGF was also combined with an empirical force field. The resulting DDGF + FF model was able to maintain the superior diversity (as reflected by the Recall metrics) that ML models are known for, while also achieving a significant boost in accuracy (as evidenced by the Precision metrics). This outcome suggests that integrating DDGF with a force field is a potent strategy, ensuring that the model retains its capacity to generate a wide range of conformations (thereby capturing the complex multimodal nature of molecular structures), while also enhancing the precision and reliability of the conformations it predicts. It’s an effective fusion of the explorative power of ML with the targeted refinement capabilities of force fields.
V-D Property Prediction
The mean absolute errors (MAE) in the context of molecular modeling are indicative of how accurately a model can predict certain properties of molecules, such as their energy levels, when compared to the ground truth obtained from experimental or highly accurate quantum chemical calculations. In this case, Table II reports the MAE for various properties of molecules that have been generated by different models. The exclusion of CVGAE from this comparison due to its subpar performance, as noted in previous work, suggests that it was significantly less accurate than other models tested. The properties in question are highly sensitive to the geometric structure of the molecules – for instance, slight changes in the distance between atoms can lead to significant differences in the predicted properties. Therefore, a model that can accurately predict these properties demonstrates that it can consistently generate molecular conformations that closely resemble true molecular structures. In summary, the low MAE reported for DDGF implies that it not only generates conformations with high diversity but also with a high degree of geometric accuracy, thus effectively predicting the properties that are closely aligned with what would be expected of the actual molecules. This reinforces the idea that DDGF is capable of capturing the complex underlying distribution of molecular structures in a way that is both diverse and precise.
TABLE II: MAE of predicted ensemble properties in eV.
Method $\bar{E}$ $E_{min}$ $\bar{\Delta\epsilon}$ $\Delta\epsilon_{min}$ $\Delta\epsilon_{max}$
RDKit 0.9233 0.6585 0.3698 0.8021 0.2359
GraphDG 9.1027 0.8882 1.7973 4.1743 0.4776
CGCF 28.9661 2.8410 2.8356 10.6361 0.5954
ConfVAE 8.2080 0.6100 1.6080 3.9111 0.2429
DDGF 0.8872 0.5962 0.3874 0.7782 0.2247
VI Conclusion
In conclusion, this study has unveiled a pioneering generative framework, DDGF, which signifies a substantial leap forward in the accurate prediction of three-dimensional molecular structures from their two-dimensional graph representations. By leveraging the principles of diffusion from non-equilibrium thermodynamics, DDGF transcends traditional bifurcated methodologies, instead fostering a coherent transformation of stochastic distributions into precise molecular conformations. Its unique bilevel optimization process, tailored to the intricacies of molecular geometry, and its commitment to preserving roto-translational invariance set new benchmarks in the field of computational chemistry. The extensive empirical evaluations presented herein underscore DDGF’s superiority over existing state-of-the-art methods, cementing its role as a critical tool for advancements in drug development and the broader understanding of molecular dynamics and interactions.